OpenStudy (anonymous):

Rewrite the expression 1/3[ln x – 2ln(1 – x)] as a single logarithm.

OpenStudy (anonymous):

1/3[ln x-2 ln(1-x)] =1/3[ln x-ln (1-x)^2] =1/3[ln{x/(1-x)^2}] =Ln[x/(1-x)^2]^1/3

OpenStudy (anonymous):

4^2=16 5^2=25 irrational numbers between 4 &5 are sqrt17,sqrt18,sqrt19,sqrt20,..........